OpenStudy (anonymous):

PLEASEEE HELP!!!! The formula to find the period of orbit of a satellite around a planet is T2=(4π2GM)r3 where r is the orbit’s mean radius, M is the mass of the planet, and G is the universal gravitational constant. If you are given all the values except r, how do you rewrite the formula to solve for r?

OpenStudy (anonymous):

you want to isolate r

OpenStudy (irishboy123):

this is the correct equation \(\large T^2=\frac{4π^2}{GM}r^3\)

OpenStudy (anonymous):

yes that's the one

OpenStudy (anonymous):

T2=(4π2GM)r3 \[( \frac{ 1 }{ 4\pi2gm})T2= r^3\] \[r=\sqrt[3]{\frac{ t2}{ 4\pi2gm }}\]

OpenStudy (anonymous):

deeez that's not one of my options... hold up... im new ill try to uploaod them for you

OpenStudy (anonymous):

upload

OpenStudy (irishboy123):

what happens if you multiply each side of the equation by GM? you can draw it if you have problems entering equations

OpenStudy (anonymous):

it gets cancelled out in the denominator and goes to the opposite side of the equation

OpenStudy (anonymous):

@Deeezzzz

OpenStudy (anonymous):

thank you

OpenStudy (irishboy123):

and then divide both sides by \(4 \pi^2\) ??